PROBLEM: (1/5 – i)^2 : PROBLEM: (1/5 – i)^2
Answer by solver91311(1877)

Use the fact that (a+b)^2=a^2+2ab+b^2 and remember that (-i)^2=(-i)(-i)=i^2=-1